# Spectrum-Aware Debiasing: A Modern Inference Framework with Applications to Principal Components Regression

### v4

@@ B.1 Proof result A: Distribution characterization
In this section, we prove Theorem 5.1 using VAMP algorithm as proof device. We define the version of VAMP algorithm we will use in Section B.1.1, prove Cauchy convergence of its iterates in Section B.1.2, and prove Theorem 5.1 in Section B.1.3. To streamline the presentation, proofs of intermediate claims are collected in Section B.3. We also assume without loss of generality that
@@
B.1.1 The oracle VAMP algorithm
@@
$F(q,x):=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}q-\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}x$ (60)
$F^{\prime}(q,x):=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}^{\prime}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}$
Note that for any fixed $x$, $F^{\prime}(q,x)$ equals to the derivative of $q\mapsto F(q,x)$ whenever the derivative exists, and at the finitely many points where $q\mapsto F(q,x)$ is not differentiable $F^{\prime}(q,x)$ equals to 00 (cf. Lemma 2.7). We also define some quantities
$\bm{\Lambda}:=\frac{\eta_{*}\left(\eta_{*}-\gamma_{*}\right)}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}-\left(\frac{\eta_{*}-\gamma_{*}}{\gamma_{*}}\right)\cdot\mathbf{I}_{p}$ (61)
$\bm{\xi}:=\mathbf{Q}\bm{\varepsilon},\quad\mathbf{e}_{b}:=\frac{\eta_{*}}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}\mathbf{D}^{\top}\bm{\xi},\quad\mathbf{e}:=\mathbf{O}^{\top}\mathbf{e}_{b}$
We note some important properties of these quantities, which are essentially consequence of Assumption 2 and (20). We defer the proof of Proposition B.1 to Section B.3.1.
Proposition B.1.
Under Assumption 24, almost surely,
@@
The following Proposition will be needed later. Its proof is deferred to Section B.3.1.
Proposition B.2.
@@
Now we state the state evolution for the VAMP algorithm. Its proof is deferred to Section B.3.1.
Proposition B.3.
Suppose Assumptions 26 hold. Further assume that the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ defined in Lemma 2.7 is non-constant. Let $\mathsf{H}=(\mathsf{B}^{\star},\mathsf{D},\mathsf{D}\Xi,\mathsf{L},\mathsf{E}_{b},\mathsf{E},\mathsf{P}_{0})$ be as defined in Proposition B.2. Set $\mathsf{X}_{1}=F(\mathsf{P}_{0},\mathsf{B}^{\star})$, set $\Delta_{1}=\mathbb{E}[\mathsf{X}_{1}^{2}]\in\mathbb{R}^{1\times 1}$, and define iteratively $\mathsf{S}_{t},\mathsf{Y}_{t},\mathsf{X}_{t+1},\Delta_{t+1}$ for $t=1,2,3,\ldots$ such that
@@
Proof of the following Corollary is deferred to Section B.3.1.
Corollary B.5.
Under Assumptions 26, almost surely as $p,n\to\infty$
@@
B.1.2 Cauchy convergence of VAMP iterates
The following Proposition is analogous to [34, Proposition 2.3] and [59, Lemma B.2.] in the context of rotationally invariant spin glass and Bayesian linear regression. However, it requires observing a simple but crucial property of the R-transform (i.e. $-zR^{\prime}(z)/R(z)<1$ for all $z$ on the domain) and its interplay with the non-expansiveness of the proximal map. We defer the proof to Section B.3.1.
Proposition B.6.
Under Assumptions 26,
@@ B.1.2 Cauchy convergence of VAMP iterates
We can then obtain the convergence of vector iterates for the oracle VAMP algorithm. We defer the proof to Section B.3.1.
Corollary B.7.
Under Assumptions 26, for $j=1,2$,
@@
B.1.3 Characterize limits of empirical distribution
Recall definition of $\mathbf{r}_{*},\mathbf{r}_{**}$ from (34). The following is a direct consequence of the Cauchy convergence of the VAMP iterates and the strong convexity in the penalized loss function. We defer the proof to Section B.3.2.
Proposition B.8.
Under Assumptions 26 with $c_{0}>0$, for $j=1,2$,
@@ B.1.3 Characterize limits of empirical distribution
Combining Proposition B.8 and Corollary B.5 yields the proof of Theorem 5.1.
Proof of Theorem 5.1.
@@
×$\qquad\qquad\times\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)$
@@
×$\qquad\qquad\times\left(\frac{1}{p}\sum_{i=1}^{p}\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)^{2}\right)^{\frac{1}{2}}$
@@
×$\qquad\qquad\times\left(3+\frac{3}{p}\left(\left\|\hat{\mathbf{x}}_{1t}\right\|_{2}^{2}+3\left\|\mathbf{r}_{1t}\right\|_{2}^{2}+2\left\|\bm{\beta}^{\star}\right\|_{2}^{2}+2\left\|\mathbf{r}_{1t}-\mathbf{r}_{*}\right\|_{2}^{2}\right)\right)^{\frac{1}{2}}$
where ()($\star$)( ⋆ ) is by Cauchy-Schwarz inequality. This, along with Proposition B.8, Assumption 3, Corollary B.5 implies that
@@
Using Corollary B.5 and Proposition A.1, we have that
@@
[…]
@@
The proof of the second result in (36) is analgous using Corollary B.5. This completes the proof. ∎
B.2 Prove result B: Consistent estimation
We prove existence and uniqueness of the solution to the adjustment equation (23) in Section B.2.1, show that the adjustment equation converges to a population limit in Section B.2.2, and prove Theorem 3.1 in Section B.2.3. To streamline the presentation, proofs of intermediate claims are collected in Section B.4.
B.2.1 Properties of the adjustment equation
Recall definition of function $g_{p}:(0,+\infty)\mapsto\mathbb{R}$ from (22). We outline in Lemma B.9 the conditions under which it is well-defined, strictly increasing and the equation
@@ B.2.1 Properties of the adjustment equation
admits a unique solution on $(0,+\infty)$. The proof is deferred to Section B.4.1.
Lemma B.9.
@@
The following assumption is made to simplify the conditions outlined in Lemma B.9.
@@
The following is a direct consequence of Lemma B.9 which in turn has Proposition 3.1 as a special case.
Proposition B.10.
@@
B.2.2 Population limit of the adjustment equation
@@
which is well-defined under Assumption 4, 6 as shown in Lemma B.11 below. We defer its proof to Section B.4.2.
Lemma B.11.
Under Assumption 4, 6, $g_{\infty}$ is well-defined on and strictly increasing on $(0,+\infty)$. The equation $g_{\infty}(\gamma)=1$ admits a unique solution $\gamma_{*}$ on $(0,+\infty)$.
Remark B.12.
We emphasize that the proof of Lemma B.11 does not require (20) admits a unique solution, only that a solution exists.
We can show that the LHS of the sample adjustment equation converges to the LHS of the population adjustment equation. We defer its proof to Section B.4.2.
Proposition B.13.
Under Assumption 26, almost surely for all sufficiently large $p$, $g_{p}$ is well-defined on $(0,+\infty)$ and for any $\gamma>0$, almost surely
@@
B.2.3 Consistent estimation of fixed points
We are now ready to prove Theorem 5.2 which shows that the quantities defined in (26) indeed converges to their population counterparts.
Proof of Theorem 5.2.
@@ B.2.3 Consistent estimation of fixed points
as a direct consequence of Proposition B.13 and that $g_{\infty}$ is strictly increasing (cf. Lemma B.11). It follows that almost surely for all $m$ sufficiently large
@@
By the assumptions and Lemma B.11, $g_{p}$ is continuous, strictly increasing and (71) admits a unique solution $\widehat{\mathsf{adj}}\left(p\right)$ for all $p$. (73) thus implies that almost surely for all $p$ sufficiently large $\left|\widehat{\mathsf{adj}}\left(p\right)-\gamma_{*}\right|<\epsilon$. This completes the proof for $\lim_{p\rightarrow\infty}\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*}$. The consistency of $\hat{\eta}_{*}$ immediately follows. To show $\hat{\bm{\beta}}^{u}(p)\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}$ almost surely as $p\to\infty$, note that $\hat{\bm{\beta}}^{u}(p)-\mathbf{r}_{*}\stackrel{{\scriptstyle W_{2}}}{{\to}}0$ by consistency of $\widehat{\mathsf{adj}}$ and $\limsup_{p\to\infty}p^{-1}\left\|\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|_{2}^{2}<+\infty$, and the claims follow from (35) and an application of Proposition A.4. A similar argument shows that $\frac{1}{p}\norm{\hat{\mathbf{{r}}}_{**}-\mathbf{r}_{**}}^{2}\rightarrow 0$ almost surely as $p\to\infty$. The consistency statements for $\hat{\sigma}^{2},\hat{\tau}_{**},\hat{\tau}_{*}$ follow from results above, (36), (20d) and Assumption 7. ∎
B.3 Supporting proofs for result A
B.3.1 Oracle VAMP proofs
Proof of Proposition B.1.
@@ B.3.1 Oracle VAMP proofs
=$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)^{2}$
@@
and in (b) we used (20c). We remark that although the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}(x)$ may not be differentiable on a finite set of points, Stein’s lemma can still be applied (cf. [79, Lemma 1]). ∎
Proof of Proposition B.2.
Note that $\bm{\xi}=\mathbf{Q}\bm{\varepsilon}\sim N(\rm{0},\mathbf{I}_{n})$. Then $\mathbf{D}^{\top}\bm{\xi}\in\mathbb{R}^{n}$ may be written as the entrywise product of $\mathbf{D}^{\top}\bm{1}_{n\times 1}\in\mathbb{R}^{p}$ and a vector $\bar{\bm{\xi}}\sim N(\rm{0},\mathbf{I}_{p})$, both when $p\geq n$ and when $n\leq p$. The almost-sure convergence $H\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{H}$ is then a straightforward consequence of Propositions A.2, A.3, and A.6, where all random variables of $\mathsf{H}$ have finite moments of all orders under Assumptions 2 and 3. The identities $\kappa_{*}=\mathbb{E}\mathsf{L}^{2}$ and $b_{*}=\mathbb{E}\mathsf{E}_{b}^{2}$ follows from definitions of $\kappa_{*},b_{*}$ in Proposition B.1. ∎
Proof of Proposition B.3.
@@
We take as base case $t=0$, where the first two statements are vacuous, and the third statement requires $(\mathbf{H},{\mathbf{x}}^{1})\stackrel{{\scriptstyle W_{2}}}{{\to}}(\mathsf{H},\mathsf{X}_{1})$ almost surely as $p\to\infty$. Recall that ${\mathbf{x}}^{1}=F(p^{0},\bm{\beta}^{\star})$, and that $F(p,\beta)$ is Lipschitz by Proposition Proposition B.1. Then this third statement follows from Propositions B.2 and A.3.
@@
So almost surely for sufficiently large $p$, conditional on $(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t})$, the law of $\mathbf{s}^{t+1}$ is given by its law conditioned on $\mathbf{U}=\mathbf{O}V$, which is (see [33, Lemma B.2])
@@
Proof of Corollary B.5.
(65) is a direct consequence of Proposition B.3, (63), Proposition A.3, Proposition A.4, (59a) and the fact that proximal map is 1-Lipschitz. To see the first result in (66), note that
@@
Using (77), (78) above, Assumption 2,3, and Proposition B.3, we obtain that almost surely as $p\to\infty$
@@
Proof of Proposition B.6.
Recall that $\delta_{tt}=\delta_{*}$ for all $t\geq 1$ from Theorem B.3. Then $\delta_{st}=\mathbb{E}\left[\mathsf{X}_{s}\mathsf{X}_{t}\right]\leq\sqrt{\mathbb{E}\left[\mathsf{X}_{s}^{2}\right]\mathbb{E}\left[\mathsf{X}_{t}^{2}\right]}=\delta_{*}$ for all $s,t\geq 1$. For $s=1$ and any $t\geq 2$, observe also that
@@
Then $g^{\prime}(\delta)=\kappa_{*}\mathbb{E}\left[\mathbb{E}[F^{\prime}(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star})\mid\mathsf{G},\mathsf{B}^{\star}]^{2}\right]\geq 0$, and a similar argument shows $g^{\prime\prime}(\delta)\geq 0$. Observe that at $\delta=\delta_{*}$, we have $\mathsf{P}_{\delta_{*}}^{\prime}=\mathsf{P}_{\delta_{*}}^{\prime\prime}=\sqrt{\sigma_{*}^{2}+b_{*}}\cdot\mathsf{G}=\sqrt{\tau_{*}}\mathsf{G}$ which is equal in law to $\mathsf{P}\sim N(\rm{0},\tau_{*})$. Then $g(\delta_{*})=\mathbb{E}[F(\mathsf{P},\mathsf{B}^{\star})^{2}]=\delta_{*}$ by Proposition B.1. So $g:[0,\delta_{*}]\to[0,\delta_{*}]$ is a non-negative, increasing, convex function with a fixed point at $\delta_{*}$. We claim that

### v6

@@ C.3 Proof result A: Distribution characterization
In this section, we prove Theorem C.1 using VAMP algorithm as proof device. We define the version of VAMP algorithm we will use in Section C.3.1, prove Cauchy convergence of its iterates in Section C.3.2, and prove Theorem C.1 in Section C.3.3. To streamline the presentation, proofs of intermediate claims are collected in Section C.5. We also assume without loss of generality that
@@
C.3.1 The oracle VAMP algorithm
@@
Remark C.2.
Note that the above definition assumes existence of fixed point $\eta_{*},\gamma_{*},\tau_{*},\tau_{**}\in(0,+\infty)$, i.e. Assumption 5. We however do not require the fixed point to be unique. Our proof may proceed by defining the oracle VAMP algorithm above with respect to any one of the fixed points.
@@
$F(q,x)=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}q-\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}x$ (73)
$F^{\prime}(q,x)=\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\operatorname{Prox}^{\prime}_{\gamma_{*}^{-1}h}(q+x)-\frac{\gamma_{*}}{\eta_{*}-\gamma_{*}}$
Note that for any fixed $x$, $F^{\prime}(q,x)$ equals to the derivative of $q\mapsto F(q,x)$ whenever the derivative exists, and at the finitely many points where $q\mapsto F(q,x)$ is not differentiable $F^{\prime}(q,x)$ equals to 0 (cf. Lemma 2.8). We also define some quantities
$\bm{\Lambda}=\frac{\eta_{*}\left(\eta_{*}-\gamma_{*}\right)}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}-\left(\frac{\eta_{*}-\gamma_{*}}{\gamma_{*}}\right)\cdot\mathbf{I}_{p}$ (74)
$\bm{\xi}=\mathbf{Q}\bm{\varepsilon},\quad\mathbf{e}_{b}=\frac{\eta_{*}}{\gamma_{*}}\left(\mathbf{D}^{\top}\mathbf{D}+\left(\eta_{*}-\gamma_{*}\right)\mathbf{I}_{p}\right)^{-1}\mathbf{D}^{\top}\bm{\xi},\quad\mathbf{e}=\mathbf{O}^{\top}\mathbf{e}_{b}$
We note some important properties of these quantities, which are essentially consequence of Assumption 1 and (20). We defer the proof of Proposition C.3 to Section C.5.1.
Proposition C.3.
Under Assumption 13 and 5, almost surely,
@@
The following Proposition will be needed later. Its proof is deferred to Section C.5.1.
Proposition C.4.
@@
Now we state the state evolution for the VAMP algorithm. Its proof is deferred to Section C.5.1.
Proposition C.5.
Suppose Assumption 13 and 5 hold. Further assume that the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}^{\prime}(x)$ defined in Lemma 2.8 is non-constant. Let $\mathsf{H}=(\mathsf{B}^{\star},\mathsf{D},\mathsf{D}\Xi,\mathsf{L},\mathsf{E}_{b},\mathsf{E},\mathsf{P}_{0})$ be as defined in Proposition C.4. Set $\mathsf{X}_{1}=F(\mathsf{P}_{0},\mathsf{B}^{\star})$, set $\Delta_{1}=\mathbb{E}[\mathsf{X}_{1}^{2}]\in\mathbb{R}^{1\times 1}$, and define iteratively $\mathsf{S}_{t},\mathsf{Y}_{t},\mathsf{X}_{t+1},\Delta_{t+1}$ for $t=1,2,3,\ldots$ such that
@@
Proof of the following Corollary is deferred to Section C.5.1.
Corollary C.7.
Under Assumption 13 and 5, almost surely as $p,n\to\infty$
@@
C.3.2 Cauchy convergence of VAMP iterates
The following Proposition is analogous to [41, Proposition 2.3] and [69, Lemma B.2.] in the context of rotationally invariant spin glass and Bayesian linear regression. However, it requires observing a simple but crucial property of the R-transform (i.e. $zR^{\prime}(z)/R(z)<1$ for all $z$ on the domain) and its interplay with the non-expansiveness of the proximal map. We defer the proof to Section C.5.1.
Proposition C.8.
Under Assumption 13 and 5,
@@ C.3.2 Cauchy convergence of VAMP iterates
We can then obtain the convergence of vector iterates for the oracle VAMP algorithm. We defer the proof to Section C.5.1.
Corollary C.9.
Under Assumption 13 and 5, for $j=1,2$,
@@
C.3.3 Characterize limits of empirical distribution
Recall definition of $\mathbf{r}_{*},\mathbf{r}_{**}$ from (61). The following is a direct consequence of the Cauchy convergence of the VAMP iterates and the strong convexity in the penalized loss function. We defer the proof to Section C.5.2.
Proposition C.10.
Under Assumptions 15, for $j=1,2$,
$\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\hat{\bm{\beta}}-\hat{\mathbf{x}}_{jt}\right\|_{2}^{2}=\lim_{t\rightarrow\infty}\lim_{p\rightarrow\infty}\frac{1}{p}\left\|\mathbf{r}_{jt}-\mathbf{r}_{j*}\right\|_{2}^{2}=0.$ (81)
@@ C.3.3 Characterize limits of empirical distribution
Combining Proposition C.10 and Corollary C.7 yields the proof of Theorem C.1.
Proof of Theorem C.1.
@@
$\qquad\qquad\times\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)$
@@
$\qquad\qquad\times\left(\frac{1}{p}\sum_{i=1}^{p}\left(1+\sqrt{\hat{x}_{1t,i}^{2}+r_{1t,i}^{2}+\beta_{i}^{\star 2}}+\sqrt{\hat{{\beta}}_{i}^{2}+r_{*,i}^{2}+\beta_{i}^{\star 2}}\right)^{2}\right)^{\frac{1}{2}}$
@@
$\qquad\qquad\times\left(3+\frac{3}{p}\left(\left\|\hat{\mathbf{x}}_{1t}\right\|_{2}^{2}+3\left\|\mathbf{r}_{1t}\right\|_{2}^{2}+2\left\|\bm{\beta}^{\star}\right\|_{2}^{2}+2\left\|\mathbf{r}_{1t}-\mathbf{r}_{*}\right\|_{2}^{2}\right)\right)^{\frac{1}{2}}$
where $(\star)$ is by Cauchy-Schwarz inequality. This, along with Proposition C.10, Assumption 2, Corollary C.7 implies that
@@
Using Corollary C.7 and Proposition A.1, we have that
@@
$\begin{aligned} &\left|\mathbb{E}\psi\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right)-\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{{\beta}}_{i},r_{*,i},{\beta}^{\star}_{i}\right)\right|\\ &\leq\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{x}_{1t,i},r_{1t,i},{\beta}^{\star}_{i}\right)-\mathbb{E}\psi\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right),\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star},\mathsf{B}^{\star}\right)\right|\\ &\qquad+\left|\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{x}_{1t,i},r_{1t,i},{\beta}^{\star}_{i}\right)-\frac{1}{p}\sum_{i=1}^{p}\psi\left(\hat{{\beta}}_{i},r_{*,i},{\beta}^{\star}_{i}\right)\right|\end{aligned}.$
@@
The proof of the second result in (141) is analgous using Corollary C.7. This completes the proof. ∎
C.4 Prove result B: Consistent estimation
We prove existence and uniqueness of the solution to the adjustment equation (25) in Section C.4.1, show that the adjustment equation converges to a population limit in Section C.4.2, and prove Theorem 3.1 in Section C.4.3. To streamline the presentation, proofs of intermediate claims are collected in Section C.6.
C.4.1 Properties of the adjustment equation
Recall definition of function $g_{p}:(0,+\infty)\mapsto\mathbb{R}$ from (24). We outline in Lemma C.11 the conditions under which it is well-defined, strictly increasing and the equation
@@ C.4.1 Properties of the adjustment equation
admits a unique solution on $(0,+\infty)$. The proof is deferred to Section C.6.1.
Lemma C.11.
@@
The following assumption is made to simplify the conditions outlined in Lemma C.11.
@@
The following is a direct consequence of Lemma C.11 which in turn has Proposition 3.1 as a special case.
Proposition C.12.
@@
C.4.2 Population limit of the adjustment equation
@@
which is well-defined under Assumption 3, 5 as shown in Lemma C.13 below. We defer its proof to Section C.6.2.
Lemma C.13.
Under Assumption 3, 5, $g_{\infty}$ is well-defined on and strictly increasing on $(0,+\infty)$. The equation $g_{\infty}(\gamma)=1$ admits a unique solution $\gamma_{*}$ on $(0,+\infty)$.
Remark C.14.
We emphasize that the proof of Lemma C.13 does not require (20) admits a unique solution, only that a solution exists.
We can show that the LHS of the sample adjustment equation converges to the LHS of the population adjustment equation. We defer its proof to Section C.6.2.
Proposition C.15.
Under Assumption 15, almost surely for all sufficiently large $p$, $g_{p}$ is well-defined and strictly increasing on $(0,+\infty)$ $g_{p}$ and equation (85) admits a unique solution on $(0,+\infty)$. Furthermore, for any $\gamma>0$, almost surely,
@@
C.4.3 Consistent estimation of fixed points
We are now ready to prove Theorem C.2 which shows that the quantities defined in (28) indeed converges to their population counterparts.
Proof of Theorem C.2.
@@ C.4.3 Consistent estimation of fixed points
as a direct consequence of Proposition C.15 and that $g_{\infty}$ is strictly increasing (cf. Lemma C.13). It follows that almost surely for all $p$ sufficiently large
@@
Since $g_{p}$ is increasing and continuous almost surely for all sufficiently large $p$, (87) implies that almost surely for all $p$ sufficiently large $\left|\widehat{\mathsf{adj}}\left(p\right)-\gamma_{*}\right|<\epsilon$. This completes the proof for $\lim_{p\rightarrow\infty}\widehat{\mathsf{adj}}\left(p\right)\rightarrow\gamma_{*}$. The consistency of $\hat{\eta}_{*}$ immediately follows. To show $\hat{\bm{\beta}}^{u}(p)\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{B}^{\star}+\sqrt{\tau_{*}}\mathsf{Z}$ almost surely as $p\to\infty$, note that $\hat{\bm{\beta}}^{u}(p)-\mathbf{r}_{*}\stackrel{{\scriptstyle W_{2}}}{{\to}}0$ by consistency of $\widehat{\mathsf{adj}}$ and $\limsup_{p\to\infty}p^{-1}\left\|\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}\hat{\bm{\beta}})\right\|_{2}^{2}<+\infty$, and the claims follow from (140) and an application of Proposition A.4. A similar argument shows that $\frac{1}{p}\norm{\hat{\mathbf{{r}}}_{**}-\mathbf{r}_{**}}^{2}\rightarrow 0$ almost surely as $p\to\infty$. The consistency statements for $\hat{\sigma}^{2},\hat{\tau}_{**},\hat{\tau}_{*}$ follow from results above, (141), (20d) and Assumption 7. ∎
C.5 Supporting proofs for result A
C.5.1 Oracle VAMP proofs
Proof of Proposition C.3.
@@ C.5.1 Oracle VAMP proofs
$\stackrel{{\scriptstyle(a)}}{{=}}\left(\frac{\eta_{*}}{\eta_{*}-\gamma_{*}}\right)^{2}\mathbb{E}\left(\operatorname{Prox}_{\gamma_{*}^{-1}h}\left(\sqrt{\tau_{*}}\mathsf{Z}+\mathsf{B}^{\star}\right)-\mathsf{B}^{\star}\right)^{2}$
@@
and in (b) we used (20c). We remark that although the function $x\mapsto\operatorname{Prox}_{\gamma_{*}^{-1}h}(x)$ may not be differentiable on a finite set of points, Stein’s lemma can still be applied (cf. [94, Lemma 1]). ∎
Proof of Proposition C.4.
Note that $\bm{\xi}=\mathbf{Q}\bm{\varepsilon}\sim N(\rm{0},\mathbf{I}_{n})$. Then $\mathbf{D}^{\top}\bm{\xi}\in\mathbb{R}^{n}$ may be written as the entrywise product of $\mathbf{D}^{\top}\bm{1}_{n\times 1}\in\mathbb{R}^{p}$ and a vector $\bar{\bm{\xi}}\sim N(\rm{0},\mathbf{I}_{p})$, both when $p\geq n$ and when $n\leq p$. The almost-sure convergence $H\stackrel{{\scriptstyle W_{2}}}{{\to}}\mathsf{H}$ is then a straightforward consequence of Propositions A.2, A.3, and A.6, where all random variables of $\mathsf{H}$ have finite moments of all orders under Assumptions 1 and 2. The identities $\kappa_{*}=\mathbb{E}\mathsf{L}^{2}$ and $b_{*}=\mathbb{E}\mathsf{E}_{b}^{2}$ follows from definitions of $\kappa_{*},b_{*}$ in Proposition C.3. ∎
Proof of Proposition C.5.
@@
We take as base case $t=0$, where the first two statements are vacuous, and the third statement requires $(\mathbf{H},{\mathbf{x}}^{1})\stackrel{{\scriptstyle W_{2}}}{{\to}}(\mathsf{H},\mathsf{X}_{1})$ almost surely as $p\to\infty$. Recall that ${\mathbf{x}}^{1}=F(\mathbf{p}^{0},\bm{\beta}^{\star})$, and that $F(p,\beta)$ is Lipschitz by Proposition Proposition C.3. Then this third statement follows from Propositions C.4 and A.3.
@@
So almost surely for sufficiently large $p$, conditional on $(\mathbf{H},\mathbf{X}_{t+1},\mathbf{S}_{t},\mathbf{Y}_{t})$, the law of $\mathbf{s}^{t+1}$ is given by its law conditioned on $\mathbf{U}=\mathbf{O}V$, which is (see [40, Lemma B.2])
@@
Proof of Corollary C.7.
(78) is a direct consequence of Proposition C.5, (76), Proposition A.3, Proposition A.4, (72a) and the fact that proximal map is 1-Lipschitz. To see the first result in (79), note that
@@
Using (91), (92) above, Assumption 1,2, and Proposition C.5, we obtain that almost surely as $p\to\infty$
@@
Proof of Proposition C.8.
Recall that $\delta_{tt}=\delta_{*}$ for all $t\geq 1$ from Theorem C.5. Then $\delta_{st}=\mathbb{E}\left[\mathsf{X}_{s}\mathsf{X}_{t}\right]\leq\sqrt{\mathbb{E}\left[\mathsf{X}_{s}^{2}\right]\mathbb{E}\left[\mathsf{X}_{t}^{2}\right]}=\delta_{*}$ for all $s,t\geq 1$. For $s=1$ and any $t\geq 2$, observe also that
@@
Then $g^{\prime}(\delta)=\kappa_{*}\mathbb{E}\left[\mathbb{E}[F^{\prime}(\mathsf{P}_{\delta}^{\prime},\mathsf{B}^{\star})\mid\mathsf{G},\mathsf{B}^{\star}]^{2}\right]\geq 0$, and a similar argument shows $g^{\prime\prime}(\delta)\geq 0$. Observe that at $\delta=\delta_{*}$, we have $\mathsf{P}_{\delta_{*}}^{\prime}=\mathsf{P}_{\delta_{*}}^{\prime\prime}=\sqrt{\sigma_{*}^{2}+b_{*}}\cdot\mathsf{G}=\sqrt{\tau_{*}}\mathsf{G}$ which is equal in law to $\mathsf{P}\sim N(\rm{0},\tau_{*})$. Then $g(\delta_{*})=\mathbb{E}[F(\mathsf{P},\mathsf{B}^{\star})^{2}]=\delta_{*}$ by Proposition C.3. So $g:[0,\delta_{*}]\to[0,\delta_{*}]$ is a non-negative, increasing, convex function with a fixed point at $\delta_{*}$. We claim that